OpenStudy (anonymous):

5^99 divided by 13 what is the reminder ?

mathslover (mathslover):

Before I go through the solution, let me ask first, are you aware about modulo or modulus method?

OpenStudy (anonymous):

nope can you answer this on the basis of binomial theorems?

OpenStudy (anonymous):

.718420321

OpenStudy (anonymous):

8

mathslover (mathslover):

Well, let me try it first, it has been a while I have tried this method. Give me some minutes

OpenStudy (anonymous):

7 or 8 depending on to what place you have toround

OpenStudy (anonymous):

@OwlGirl212 how did u get that?

OpenStudy (anonymous):

Working mod 13, note that 5^(13-1) = 1 (mod 13). So, 5^99 = 5^(12 * 8 + 3) .............= (5^12)^8 * 5^3 .............= 1^8 * 125 .............= 8 (mod 13). Since 8 is between 0 and 12 inclusive, the remainder in question equals 8.

mathslover (mathslover):

@OwlGirl212 - The solution you posted doesn't belong to you! It is considered as Plagiarism. Quoting some other person's work without stating the owner or writer of that part is considered as Plagiarism. Please prefer to post your own solution in future. If, due to any reasons you have to copy-paste a solution from other resource, then do mention the link Here is from where you copied it : https://in.answers.yahoo.com/question/index?qid=20120521102952AA9NyEn

mathslover (mathslover):

@ganeshie8 - Can you help him? I am not sure whether I will be able to help him ... I have to leave now. Sorry!

OpenStudy (anonymous):

someone please help me ..

ganeshie8 (ganeshie8):

\(\large (x+y)^n = \sum \limits_{k=0}^n \binom{n}{k} x^{n-k}y^k\)

ganeshie8 (ganeshie8):

\(\large 5^{99} = 5\times 5^{98} = 5\times (5^2)^{49} = (26-1)^{49} \)

ganeshie8 (ganeshie8):

Notice that 26 leaves 0 as remainder when divided by 13.

ganeshie8 (ganeshie8):

think wat terms u wud get when u expand above^

ganeshie8 (ganeshie8):

\(\large (26-1)^{49} = \sum \limits_{k=0}^{49} \binom{49}{k} 26^{49-k}(-1)^k \)

ganeshie8 (ganeshie8):

\(5^{99} = 5\times 5^{98} = 5\times (5^2)^{49} \\ = 5 \times (26-1)^{49} = 5 \times \left( 26(somthing) -1\right)\)

ganeshie8 (ganeshie8):

divide it by 13, whats the remaider ?

OpenStudy (anonymous):

divide what?

ganeshie8 (ganeshie8):

\(\large \frac{5 \times \left( 26(somthing) -1\right)}{13} \)

OpenStudy (anonymous):

but how to divide that eq?

ganeshie8 (ganeshie8):

easy, keep in mind what you're after

ganeshie8 (ganeshie8):

\(\large \frac{5 \times \left( 26(somthing) -1\right)}{13} \) \(\large \frac{5 \times 26(somthing) -5}{13} \)

ganeshie8 (ganeshie8):

since 13 divides 26 fully, any multiple of 26 wud leave 0 as remainder

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

\(\large \frac{5 \times 26(somthing) -5}{13} \) \(\large \frac{5 \times 26(somthing) }{13} - \large \frac{5}{13} \) \(\large 0 - \large \frac{5}{13} \)

ganeshie8 (ganeshie8):

^thats the remainder

ganeshie8 (ganeshie8):

-5 or -5+13 = 8 both are same

OpenStudy (anonymous):

how did that +13 come?

ganeshie8 (ganeshie8):

you can add 13's when u divide by 13

ganeshie8 (ganeshie8):

adding 13's will not affect ur remainder

ganeshie8 (ganeshie8):

13 is same as 26 is same as 39 is same as -13 when u talk about remainders when divided by 13

OpenStudy (anonymous):

so if i get a negative remainder i need to add it with the divisor??

OpenStudy (ikram002p):

5^12=1 mod 13 5^(12*8)=1^8 mod 13 5^96=1 mod 13 5^3=8 mod 13 5^99=8 mod 13

ganeshie8 (ganeshie8):

if u want a positive answer, yes

ganeshie8 (ganeshie8):

saying the remainder is "-5" is perfectly fine as well

ganeshie8 (ganeshie8):

but in general ppl like positive numbers as remainders, so... add 13

OpenStudy (cwrw238):

@ganeshie8 - this is probably a silly question but I can't see how you got 5*(26(something - 1)

OpenStudy (cwrw238):

just that last step

ganeshie8 (ganeshie8):

oh.. sorry i dint show some steps in between... one sec... :)

ganeshie8 (ganeshie8):

\(\large 5^{99} = 5\times (26-1)^{49} = 5\sum \limits_{k=0}^{49} \binom{49}{k} 26^{49-k}(-1)^k \) \(= 5 \left[ \binom{49}{0} 26^{49-0}(-1)^0 + \binom{49}{1} 26^{49-1}(-1)^1 + ... +\binom{49}{49} 26^{49-49}(-1)^{49} \right] \) \( =5 \left[26(somthing) + \binom{49}{0} 26^{49-49}(-1)^{49} \right] \)

ganeshie8 (ganeshie8):

in the binomial expansion, all terms except the last term has a factor 26 in common... so we can pull it out

ganeshie8 (ganeshie8):

corrected the typo : \(\large 5^{99} = 5\times (26-1)^{49} = 5\sum \limits_{k=0}^{49} \binom{49}{k} 26^{49-k}(-1)^k \) \(= 5 \left[ \binom{49}{0} 26^{49-0}(-1)^0 + \binom{49}{1} 26^{49-1}(-1)^1 + ... +\binom{49}{49} 26^{49-49}(-1)^{49} \right] \) \( =5 \left[26(somthing) + \binom{49}{\color{red}{49}} 26^{49-49}(-1)^{49} \right] \)

OpenStudy (cwrw238):

ah i see it now thanks

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

thank u so much guys